A mass of ideal gas of volume 400 cm3 at a temperature of 27 ˚C expands adiabatically
until its volume is 500 cm3. Calculate the new temperature. The gas is then compressed
isothermally until its pressure returns to the original value. Calculate the final volume of
the gas. Assume γ = 1.40.
1
Expert's answer
2020-07-02T17:13:48-0400

Solution.

"V_1=400cm^3;"

"T_1=27^oC=300K;"

"V_2=500cm^3;"

"\gamma=1.4;"

1) "\dfrac{T_2}{T_1}=(\dfrac{V_1}{V_2})^{\gamma-1};"

"T_2=T_1(\dfrac{V_1}{V_2})^{\gamma-1};"

"T_2=300K(\dfrac{400cm^3}{500cm^3})^{0.4}=274K;"


"P_1V_1^\gamma=P_2V_2^\gamma; \dfrac{P_1}{P_2}=(\dfrac{V_2}{V_1})^\gamma;"


"\dfrac{P_1}{P_2}=(\dfrac{500cm^3}{400cm^3})^{1.4}=1.37;"


2)"P_1V_1=P_2V_2 -" isothermal process;


"\dfrac{P_1}{P_2}=\dfrac{V_2}{V_1} \implies V_2=\dfrac{P_1}{P_2}V_1;"


"V_2=\dfrac{1}{1.37}\sdot500cm^3=365cm^3;"


Answer:1) "T_2=274K;"

2) "V_2=365cm^3."